I need HELP!
1. 6 = 2y - y + 4 4. 3(7 - 11) = y - 5

2. 4 - 9 = a - 1 + 14

3. -23 + 8 + x = -2 + 13

Answer by bonster(299)

1. 6 = 2y - y + 4
6=y+4
Subtract 4 from both sides
(6)-4=(y+4)-4
2=y

2. 4 - 9 = a - 1 + 14
-5=a+13
Subtract 13 from both sides
(-5)-13=(a+13)-13
-18=a

3. -23 + 8 + x = -2 + 13
-15+x=11
add 15 to both sides
(-15+x)+15=(11)+15
x=26

4. 3(7 - 11) = y - 5
3(-4)=y-5
-12=y-5
add 5 to both sides
(-12)+5=(y-5)+5
-7=y